### v1

# Crepe: A Mobile Screen Data Collector Using Graph Query

Abstract.
Collecting mobile datasets remains challenging for academic researchers due to limited data access and technical barriers. Commercial organizations often possess exclusive access to mobile data, leading to a “data monopoly” that restricts the independence of academic research. Existing open-source mobile data collection frameworks primarily focus on mobile sensing data rather than screen content, which is crucial for various research studies. We present Crepe, a no-code Android app that enables researchers to collect information displayed on screen through simple demonstrations of target data. Crepe utilizes a novel Graph Query technique which augments the structures of mobile UI screens to support flexible identification, location, and collection of specific data pieces. The tool emphasizes participants’ privacy and agency by providing full transparency over collected data and allowing easy opt-out. We designed and built Crepe for research purposes only and in scenarios where researchers obtain explicit consent from participants. Code for Crepe will be open-sourced to support future academic research data collection.
journalyear: 2023copyright: rightsretainedjournal: PACMHCIjournalyear: 2024journalvolume: 6journalnumber: UISTarticle: 1publicationmonth: 11doi: XXXXXXX.XXXXXXX
1. Introduction
Datasets collected from apps running on mobile devices are valuable resources in Human Computer Interaction (HCI) and social sciences (Raento et al., 2009), but collecting mobile app datasets has remained challenging for academic researchers. In practice, access to data siloed in mobile apps (Chandramouli et al., 2015) is often exclusively possessed by the smartphone platforms (e.g. iOS, Google Android) or individual products’ developers (Generosi et al., 2020). Academic researchers have to either use public developer APIs (Trupthi et al., 2017) or directly collaborate with products’ creators (Kramer et al., 2014) to obtain mobile datasets. However, many researchers have limited resources and technical expertise to build their own data collectors. Meanwhile, the availability of developer APIs are subject to corporations’ discretionary policies (Reddit, 2023; Barnes, 2023). Also, when researchers collaborate with commercial organizations to access mobile data (Kramer et al., 2014; Deka et al., 2017), they often need to align with the organizations’ business goals and public images, restricting the independence of academic research (Kramer et al., 2014). These phenomena are recognized as commercial organizations’ “data monopoly” in previous work (Jia-Jun Li et al., 2022) and pose challenges for empirical research related to mobile devices.
Many open-source mobile data collection frameworks have been created to support academics; however, most mature solutions focus on mobile sensing data (Ferreira et al., 2015; Aharony et al., 2011; Cardone et al., 2013; Tangmunarunkit et al., 2015; Schueller et al., 2014; Ferreira et al., 2015; Andone et al., 2016; Cao et al., 2015) instead of screen information collection. Screen information i.e., the various data displayed on mobile user interface (UI) screens, is an important data source to enable various research studies (Thomas et al., 2020; Robinson et al., 2017; Supanitayanon et al., 2020; Hill and Monroy-Hernández, 2017). A major technical challenge is the difficulty to reliably and automatically identify which user interface screens contain the target data, and then locate and collect it. As a result, many mobile screen collectors require users to actively input the data or upload screenshots containing the target data (Calacci and Pentland, 2022). Some collectors continuously record user screens to get the target information, which is less efficient and poses privacy concerns (Krieter, 2019).
In this work, our goal is to build a low-code tool to support the collection of screen information and user interaction data on mobile devices. To tackle the above challenges in mobile data collection, we designed and implemented a novel query language, Graph Query, which augments mobile UI screen structures to support the flexible identification, location, and collection of specific data pieces. On top of Graph Query, we designed and built an Android collector app named Crepe [^1] Crepe stands for “Collector for Research Experiments of Participant Experiences”.. Crepe supports researchers to collect screen data through simple, no-code demonstrations of the target data to collect, as well as easy, scalable sharing of collectors with participants.
We designed Crepe with a special focus on data collection participants’ privacy. The participants can download Crepe and easily join a data collection study by entering a collector ID shared by a researcher. Users will have full transparency over what data is collected from their phones and can leave any data collection at any time. We intend Crepe to be used for research purposes only, and require researchers to obtain explicit consent from participants and fully disclose the data being collected to participants.
In all, our contributions include:
  1. (1)
    A novel query language to effectively identify, locate, and collect target data on UI screens;
  2. (2)
    The first Android data collector app, Crepe, to support flexible, customizable, low-code UI screen data collection using a programming by demonstration (PBD) paradigm (Cypher and Halbert, 1993);
  3. (3)
    A user study with 7 participants in a realistic data collection task that uncovered the real-world performance, limitations, and future development directions of Crepe.
2. Background and Related Work
2.1. Understanding Human Activities via Mobile Data
Mobile devices, especially smartphones, are now ubiquitous in everyday lives. Mobile applications harness various information about user social activities, which are embedded in users’ interactions, app usage patterns, and preferences. As a result, researchers in HCI and social science have been leveraging rich data gained from mobile apps to better improve user experiences (Deka et al., 2017; Wang et al., 2023), uncover societal trends (Vieweg et al., 2010; Starbird et al., 2014; Starbird and Palen, 2010), and understanding user behaviors and attitudes (Fakhradyan, 2021; De et al., 2017; Balaban et al., 2022; Kramer et al., 2014).
Screen data from mobile phones, i.e. information displayed on mobile screens, are particularly valuable for understanding user behavior and preferences. Screen data capture visual information and layout of the apps, providing insights into how users navigate and engage with different features. By analyzing this rich, granular data, researchers can gain a deeper understanding of user needs, preferences, and pain points, which can inform the design of more user-friendly and engaging mobile experiences (Ballard, 2007).
2.1.1. Utilizing Behavioral Data Stream for Analysis
Researchers from a diverse background, such as social science or business, mine mobile data for behavioral patterns and insights. Experience Sampling Method (ESM) is a research methodology to obtain in situ data for constructing an understanding of subjects’ daily behaviors, feelings, and thoughts of participants (van Berkel et al., 2017; Yue et al., 2014). In the survey conducted by Berkel et al., smartphone-based ESM benefits researchers in improving data quality through validation, collecting rich multimodal data for context reconstruction, and enabling real-time data collection and analysis (van Berkel et al., 2017). ESM can provide social scientists with combined human and sensor data in addition to background logging (Raento et al., 2009). Aside from sensor data, ESM questionnaires can also capture external aspects of experience, such as time, place, activities, and companion (Hektner et al., 2007). For instance, Yue et al. adopted ESM to study the function of photos. They found that photos can be used to trigger memory during follow-up interviews and as a beneficial component in data analysis (Yue et al., 2014). In this work, our goal is to create a low-code mobile UI screen data collection tool. Our app Crepe can facilitate academic researchers to instrument screen data collections without the need to extensively develop customized data collectors, or have public access to data APIs.
2.2. Computational Understanding of Mobile UI Structures and Data
Technical researchers have developed intelligent systems that achieve computational understanding of mobile user interfaces. Sugilite, a mobile PBD system that the execution of various tasks through user-driven multimodal direct manipulations, parses UI elements and layout information to understand the intentions contained in user activity (Li et al., 2017). The Sugilite system leverages the accessibility API of Android to encapsulate high-quality relevant behavioral data on mobile devices. UI understanding and summarization is another topic that HCI researchers are interested in. An accurate and succinct description of UI semantics can facilitate seamless interactions as well as bridge the gap between language and user interface. Prior research such as Screen2Vec (Li et al., 2021) and Screen2Words (Wang et al., 2021), adopted Machine Learning (ML) techniques to support UI design. More recent approaches have also leveraged the advance of Large Language Models (LLMs) and large multimodal models to enhance conversational interactions on mobile UI (Wang et al., 2023; Yan et al., 2023). In this work, with our proposed novel Graph Query, we aim to create a deterministic and reliable query language that can reliably and deterministically identify and locate a piece of target information on screen. This compliments existing research on mobile UI understanding in addition to existing deep-learning-based mechanisms.
2.3. State-of-the-art Mobile Data Collection Tools
A variety of data collection tools and frameworks have been proposed to support academic data collection of various types of data. Aware is a mobile instrumental tool designed for collecting usage context through sensors on mobile devices (Ferreira et al., 2015). Rather than being a simple mobile data collector, Aware creates a collaborative framework for researchers to share context data with the community. Purple Robot is an extensible and modular development platform that supports behavioral and clinical intervention (Schueller et al., 2014). The tool enables Behavioral Intervention Technology (BIT) stakeholders to create mobile apps that collect user management, content authorship, and content delivery data, through which physician-scientists can evaluate and share data resources to increase knowledge finding and generalization. ShiptCalculator is a research tool designed to track and share workers’ aggregated data about their pay to empower workers in offering awareness of wage transparency and advocacy to campaigns. In the data collection process, Workers send a screenshots of their pay history to ShiptCalculator, which parses them using OCR and stores structured data. The system then sends a validation text to workers and allows them to explore their pay details  (Calacci and Pentland, 2022). However, no current tool supports the flexible collection of data displayed on mobile UIs based on our review of previous work. Our design of Crepe in this study aims to solve this exact problem.
3. Proposed Method
Our goal for the Crepe data collector is three-fold. First, we seek to empower researchers lacking programming expertise to easily instrument data collections. Second, we prioritize the privacy and agency of individual data collection participants. Lastly, we also want to address the following observed technical challenges in reliable mobile screen data collection, which hinder even researchers with technical skills:
  1. (1)
    The difficulty in flexibly and automatically detecting the appearance of target data on the screen;
  2. (2)
    The lack of precision in reliably pinpointing the location of target data;
  3. (3)
    The limited generalizability to diverse data types, especially for dynamically evolving data.
To get around these challenges, previous collectors have asked participants to either directly input target data or upload screenshots or screen recordings containing target data  (Calacci and Pentland, 2022; Krieter, 2019). However, these practices significantly increase participants’ efforts and reduce the reliability of the collected data’s quality.
For the rest of this section, we first give an overview of the experience design of Crepe, our Android mobile data collection app. Then, we dive into the technical details of the novel Graph Query we designed and implemented as the backbone of Crepe.
3.1. The Crepe Data Collector User Experience
Figure 2. The workflow of Crepe for our main user groups: data collection researchers and participants. Researchers create a new data collector by demonstration (A) and share the collector ID with participants (B). Participants add the collector to their own devices (C), which runs in the background to collect the specified target data (D). The collected data is transmitted to a database (E) for the researcher to analyze and gain insights (F). The colors in the background indicate each user group’s experience involved in the holistic Crepe pipeline.
[…]
The figure illustrates the workflow of the Crepe system with 6 labeled steps: A) The researcher creates a new data collector by specifying the time period, app package, data field and description through a graphical interface on their device. B) The researcher shares the collector ID with participants. C) Participants add the collector to their own devices by entering the shared collector ID into the Instagram app running on their phone. The collector details like start/end time are shown. D) On the participant’s device, the Crepe app automatically identifies and locates the target UI data elements (like ”walmart”, ”deucelove”) using Graph Query, based on the UI screenshot. E) The target data elements collected by Crepe on participant devices get transmitted to a database. F) The researcher analyzes the collected data from the database to gain insights for their study. Arrows show the flow of information between the researcher device, participant devices, and the data collection database.
The Crepe data collector has two main target user groups: researchers who want to instrument a mobile data collection study, and study participants who want to contribute mobile data to ongoing data collections [^2] We intend Crepe to be used for research purposes only where participants give their explicit consent to researchers before data collection begins..
Researchers have two main tasks when using Crepe (Figure 2): creating a collector (A), and sharing collectors with study participants (B). A researcher can use Crepe’s graphical user interfaces to define the start and end dates of a collector, as well as the target app to collect data from.
The most important yet trickiest step in defining a data collector is specifying the target data on screen to collect. In Crepe, our design goal is to minimize the complexity of this specification process for researchers, by only asking them to tap on the screen data to collect in the target app, and describe their intention to collect these data (Figure 2 (A)). Crepe automatically translate these two pieces of information into a formal, executable Graph Query, which includes information about the target data’s parent app package, component type, and relations with other UI entities on screen. This Graph Query is the core of a data collector and will be used to reliably identify and precisely locate target data on the study participants’ devices. This specification process follows a programming-by-demonstration paradigm (Cypher and Halbert, 1993) and provides a simple, no-code experience for researchers to specify the target data. (The technical detail will be further explained in the following subsection on Graph Query.) After the target data are specified and descriptions are added, a collector is created. Afterward, the researcher can copy and share the unique ID of this collector with participants (B). The collector will run on the participants’ device for the specified data collection periods. Afterwards, the researcher can receive the data collected in a self-defined database (E) and analyze it to uncover research insights (F). Our initial implementation of Crepe transmits collected data to a pre-configured Google Firebase Realtime Database that encrypts data both in transit and at rest [^3] https://firebase.google.com/support/privacy. We plan to support researchers’ customized configuration of their own database servers in the future.
For study participants, they can easily add a collector to Crepe by entering the unique collector ID shared by the study’s researchers (C), using Crepe’s graphical user interfaces. After granting Crepe with the necessary Android Accessibility Service permission, the added collector(s) will start automatically running in the background of the participants’ devices (D). Guided by the graph queries associated with the participants’ added collectors, Crepe will selectively seek the appearance of target data exclusively when the designated target apps are active. Once the target data is detected, Crepe will scrape and transmit the relevant information to the researchers (E). The participant will be able to monitor the active status of the collector(s) they have joined in Crepe. To drop out of a study, participants can simply delete a collector from their Crepe app, and the collector will no longer run in the background.
Privacy considerations
Protecting the privacy and agency of study participants is critical in our design of Crepe. We ensure individual participants have the full control over their participation status and data contributions using two main measures. First, we adopt an “opt-in” participation mechanism, where participants actively add collectors to their Crepe app to join data collections. We intend this to be combined with the explicit consents that researchers must obtain before each data collection study starts. Second, each participant can opt out of any collection, at any time, by deleting the collector from their Crepe app, which will immediately terminate the collection on their mobile devices.
For full transparency, the participant can in real-time check the examples of data being collected from their phone in the app’s collector details popup. Whenever a collector is modified by the researcher (e.g. extended or shortened collection period, additional or reduced types of data being collected), all participants will be informed of the changes through push notifications. We are also continuously rolling out new privacy-centric features, such as a data visualization dashboard for participants to see their contributed data volume and content, and the ability to request all their data be deleted when they withdraw from a study.
3.2. Graph Query
3.2.1. Building blocks of graph queries: screen entities and their relations
Figure 3. Details regarding the process of query generation and execution in Crepe. Section A (1) shows a UI Snapshot with UI elements and their relations, from which potential graph queries are generated (2) by combining screen entity relations to uniquely identify the target UI element on this UI Snapshot. Section B (3) depicts the execution of a chosen graph query on a different UI Snapshot, where the target element “duecelove” is successfully located, while the query fails to find a match on another screen.
[…]
The figure consists of two main sections, A and B, presented side by side. Section A, labeled “Query Generation”, displays a UI snapshot from a mobile app with various GUI elements like buttons, images, and text. It shows the process of creating graph queries by combining the relations between these screen entities. Section B, titled “Query Execution”, demonstrates the execution of a selected graph query on different UI snapshots. The query successfully identifies and locates the target element “duecelove” on one snapshot but fails to find a match on another screen. The figure uses visual cues like arrows and graphs to illustrate the flow of query generation and execution.
We develop a novel Graph Query in Crepe for accurate and consistent extraction of various types of screen information (Figure 3). The intuition behind Graph Query is to uniquely identify and locate the element(s) that contain the target data using the UI elements’ own attributes and relations with other elements on screen. Graph Query builds on top of the XML mobile screen hierarchy, similar to the DOM tree of web HTML, that is accessible through Android Accessibility Service. This mobile screen hierarchy contains Views on the current screen (1), each of which contains rich properties including content, screen location, etc.
First, we enhance the default mobile screen hierarchy provided by Android Accessibility Service into an augmented UI Snapshot, which contains additional relations between screen entities (Figure 3, UI Snapshot section). We defined a screen entity to be either an UI element (i.e., Android View) or an attribute of the UI element, in data representations such as strings or integers. The relations between these screen entities include:
  1. (1)
    UI elements to their own Android implementation attributes:
    • e.g. hasClassName, hasScreenLocation, isEditable
  2. (2)
    UI elements to their semantic content:
    • e.g. hasText, containsEmailAddress, containsMoney, containsDate
  3. (3)
    hierarchical relations between UI elements in the XML screen hierarchy:
    • e.g. hasParent, hasChild, hasSibling
  4. (4)
    spatial relations between UI elements on screenshot images:
    • e.g. right, left, above, below, near
These relations comprehensively encapsulate the various relations and semantic contents tha are useful for identifying and locating a date on screen. With such additional information, an UI Snapshot is a collection of subject-predicate-object triples denoted as (s, o, p), where s and o are two entities and the p is the relation between s and o.
3.2.2. Automatic creation of graph queries
When a researcher specifies a piece of data to collect through the demonstration of tapping on the target data on screen (e.g. tapping on “apple” on the screenshot in Figure 3 section 1), we locate the associated UI element in the UI Snapshot, and generate a list of potential graph queries based on the target data’s relations with other screen entities in the UI Snapshot (2).
The Graph Query can be connected via 3 logical operators: conj (and), or (or), and prev (previous). We defined a set of context-free grammars, summarized in Table  1, to construct formal graph queries based on screen entity relations in UI Snapshots.
Expression Rule
E $\rightarrow$ e;
E $\rightarrow$ S;
S $\rightarrow$ (join r E);
S $\rightarrow$ (and S S);
T $\rightarrow$ (ARG_MAX r S);
T $\rightarrow$ (ARG_MIN r S);
Q $\rightarrow$ S — T;
Table 1. Context-free grammars (CFGs) for constructing Graph Query in Crepe. Q denotes the initial non-terminal symbol, terminal e denotes a GUI object entity, and terminal r denotes a relation. Other non-terminal symbols are employed for intermediary steps in the derivation process. We used this set of CFGs to automatically construct formal Graph Queries from the screen entity relations in UI Snapshots.
Specifically, Graph Queries encapsulate combinations of the screen entity relations for the target UI element and Crepe creates combinations can uniquely identify only the target UI element on screen (Figure 3 (2)). The combinations of screen entities are empirically set through experimentation. All graph queries will have hasPackageName and hasClassName to establish the context. In most cases, more than one graph query can uniquely locate the target UI element on the screen. In these scenarios, to ensure the final selected Graph Query reflects the user’s intention of collecting the target data, we used a large language model [^4] We used GPT-3.5-turbo in our implementation due to its ease of access, short inference time, and decent performance observed on our task., to select the Graph Query that best matches users’ description of target data. The description of the target data to collect was an input field in Crepe’s graphical user interface (the details of our prompt can be found in Appendix A.1). Altogether, Crepe can effectively generate graph queries that can reliably, uniquely identify and locate the target UI element on screen.
3.2.3. Graph Query execution on UI Snapshots: retrieving the target data
After a graph query is created, it can be used to collect target data on any UI Snapshot with elements sharing the same set of entity relations. In the background, Crepe runs the Graph Query every time a relevant UI snapshot is updated by changes in screen content (Figure 3 Section (B)). We also limit the Graph Query to only run in the target app package, in order to effectively reduce unnecessary battery usage of Crepe. After idenfitying and locating the target UI element on screen (3), Crepe can collect the target data to collect through the element’s attributes such as text and content description.
3.2.4. Summary
Graph Query showcases three major strengths in screen data collection. First, it provides much more flexibility in identifying and locating target UI elements through their implementation attributes, semantic content, hierarchical screen structure position, and screen location relations (3.2.1). For instance, a researcher can utilize the pre-defined hasPrice relation to collect Uber drivers’ task payment information, a common task on gig worker support platforms (Calacci and Pentland, 2022). In the future, developers can also extend existing entity relation types for more customized data collection tasks. Second, the data collection mechanism using Graph Query is fully deterministic and more reliable than alternative solutions like using Optical Character Recognition (OCR) or deep-learning based detection models. Lastly, by utilizing Android’s Accessibility Service, using graph query can easily integrate the collection of user interaction data, a feature included in Crepe yet unsupported in most vision-based data collection tools. In the future, Crepe can be extended to collect more rich media information such as screenshots under the study participants’ consent555 Link to the documentation of Android Accessibility Service’s API for taking screenshots..
3.3. Implementation
Crepe was implemented using Android, Google Firebase Realtime Database, OAuth, and OpenAI API. The app was developed in Java using Android Studio and is compatible with mobile devices running Android 9.0 or above, which in March 2024 takes up around 90% of the Android mobile phone market. Firebase Realtime Database was selected due to its high-standard data encryption both in transit and at rest, real-time data synchronization, and reliable user authentication. To implement the query language, Crepe leverages the Accessibility Service and captures Accessibility Events as a result of changes in the content of the screen. The captured Accessibility Event contains the new UI screen hierarchy [^6] https://developer.android.com/reference/android/view/accessibility/AccessibilityEvent. The code for Crepe will be open-sourced to enable easy adoption and further development within the research community.
4. Example Use Case and User Testing
Participant ID Mobile OS Device Period Battery Usage
P1 Android 14 Pixel 8 (own) 72 hours 1%
P2 Android 14 Pixel 6 (provided) 72 hours 14%
P3 Android 14 Pixel 6 (provided) 72 hours 9%
P4 Android 13, OxygenOS 13.1 OnePlus 9R (own) 24 hours 0.91%
P5 Android 12, Oxygen OS 12.1 OnePlus Nord (own) 24 hours N/A
P6 Android 14 Pixel 6 (provided) 24 hours 2%
P7 Android 9 Galaxy S10 (own) 24 hours 0%
Table 2. Mobile device usage data for study participants over 24 to 72 hour periods, including mobile OS, device model, and battery usage percentage. Three of the participants did not have an Android phone and used our provided device, with which they were asked to use Instagram following their regular habits during the study period. The participants decided the time period of the study based on their time schedule. Note that P5 reported that they could not found the battery usage of Crepe after the collection. We were not able to help them find it either, despite our best efforts.
In this section, we describe a potential use case of collecting Instagram Story advertisements to demonstrate the collection mechanism and utility of Crepe. We also deployed a user study to understand Crepe’s collection performances and limitations in the wild.
4.1. Example Use Case: Collecting Instagram Story Advertisements
Instagram Story advertisements are sponsored short image/video segments that appear between an Instagram user’s following content. They usually come with a “Sponsored” label to be distinguished from users’ regular following content. Mutiple research studies have been conducted in recent years around Instagram Story advertisements, to understand how users perceive and understand this specific form of advertising (Balaban et al., 2022; Haldborg Jørgensen et al., 2023; Fakhradyan, 2021). However, our the best of our knowledge, no academic research to date has been able to collect a dataset of Instagram story advertisements. Such data is also not available from Instagram’s developer API  [^7] https://developers.facebook.com/docs/instagram-api/.
We believe that getting access to a dataset of Instagram’s Story advertisements for individual users can enable research contributions in fields including social media research (Hanbazazh and Reeve, 2021), algorithm auditing (Ulbricht and Yeung, 2022; Eslami et al., 2015, 2016), and usable privacy (Chen et al., 2024). Specifically, we can empirically observe users’ interactions with Instagram Story ads, as well as inferring dwelling time over Story ads based on the first and last content change while a specific ad appears on screen. More generally, some broader research questions we can answer with the dataset include:
  1. (1)
    What empirical understanding of Instagram’s personalized advertisement algorithms can we obtain from quantitative data analysis?
  2. (2)
    How Instagram Story advertisements reveal the platform’s portraits of individual users’ preferences?
  3. (3)
    What are users’ folk theories of Instagram’s personalized advertisement algorithm?
4.2. User Testing
We conducted a user study to evaluate the performance and limitations of Crepe in real-world scenarios. It is important to note that the primary objective of this study was not to collect high-quality, immediately usable datasets to support future research, but rather to assess the real-world performance of our data collector and its impact on participants’ regular mobile device usage experience. To this end, the study aimed to address the following research questions:
  1. (1)
    RQ1: How does Crepe perform in a real-world data collection scenario?
  2. (2)
    RQ2: What limitations does Crepe have when collecting data in the wild?
  3. (3)
    RQ3: How does Crepe impact the regular everyday mobile phone device usage experience of participants?
Participants
We recruited seven participants through social media advertisements and word-of-mouth. The participants are regular users of Instagram and Instagram Story. Each participant decided to contribute their data from 24 hours up to 72 hours based on their preferences and time schedule. We aimed at testing Crepe on various Android versions and device platforms, while also evaluating Crepe’s performance over short and longer collection duration. Four participants used their own Android devices. Since our main goal is not to collect a dataset that can be directly used in rigorous quantitative analysis but to empirically test our collector system, we also recruited three participants who are interested but do not have an Android device (P2, P3, P6). We provided each of them with an Android phone, with which they were asked to use Instagram during the period of the study, roughly following their regular Instagram usage habits. These three participants also contributed to our study by helping us assess the performance and usability of Crepe across different usage behaviors.
Study procedure
The user study followed a three-stage procedure [^8] The study protocol was reviewed and approved by the Institutional Review Board (IRB) at our institution.:
  1. (1)
    A 30-minute introductory interview to familiarize participants with the study and set up the Crepe app on their devices;
  2. (2)
    A testing period lasting between 24 to 72 hours, based on the participants’ time schedule and preferences, during which they were encouraged to use Instagram as usual.
  3. (3)
    A 30-minute concluding interview to administer a post-study questionnaire, collect feedback on participants’ experiences, and address any questions or concerns.
Throughout the study, we maintained email communications with participants to ensure Crepe functioned normally and conducted troubleshooting sessions when necessary. Each participant was compensated with 25 USD digital gift card for the data they contributed, and 20 USD for each hour of meeting sessions they attended. The participants’ participation duration, device information, and Crepe’s battery usage result after the study ended are shown in Table 2.
Figure 4. Number of unique Instagram Story Ads collected by participants using the Crepe data collector. The top subplot displays data from Participants 1-3 over a 72-hour period, while the bottom subplot shows data from Participants 4-7 over a 24-hour period. The x-axis represents the time intervals in hours, and the y-axis indicates the number of Instagram Story Ads collected within each time interval.
[…]
The figure consists of two vertically stacked subplots, both with multi-colored bar graphs. The top subplot, titled ”Participants 1-3 (72 hours),” displays data for three participants over a 72-hour period. Each participant’s data is represented by a different color: Participant 1 in blue, Participant 2 in purple, and Participant 3 in green. The bars indicate the number of unique Instagram Story Ads collected by each participant at various time intervals, with the time in hours along the x-axis and the number of ads on the y-axis. The bottom subplot, titled ”Participants 4-7 (24 hours),” presents data for four participants over a 24-hour period. The participants’ data are represented by different colors: Participant 4 in yellow, Participant 5 in orange, Participant 6 in dark blue, and Participant 7 in light green. Similar to the top subplot, the bars show the number of unique Instagram Story Ads collected by each participant at different time intervals, with the time in hours along the x-axis and the number of ads on the y-axis. Both subplots have clearly labeled x and y axes, and a legend is provided to identify the participants by their assigned colors. The figure allows for a visual comparison of the Instagram Story Ad data collected by different participants over their respective data collection periods.
4.3. Study Results
After initial data analysis, Crepe was able to collect 358 Instagram Story advertisements from all seven participants, with an average of 51 ads per participant (Figure 4). The collected data reveals the number of unique Instagram Story Ads encountered by each participant at regular time intervals throughout their respective data collection periods. The data allow us to observe the frequency and distribution of Instagram Story Ads over time, potentially providing insights into the advertising strategies employed by businesses on the platform. We shared the data we collected from each participant with them individually upon request.
In the post-study questionnaire, which included two 5-scale questions, all participants unanimously reported that they felt minimal interruption and disturbance from Crepe while using their everyday apps and did not notice significant changes to their devices during the collection study. Here are details regarding these two questions and all participants’ response:
  • Q1: When using my everyday apps, I felt minimal interruption and disturbance from the Crepe app.
    • All participants responded 5 – Strongly agree
  • Q2: I noticed significant changes to my device during the data collection study.
    • All participants responded 1 – Strongly disagree
Based on the battery usage data collected in the post-study questionnaire, Crepe demonstrated excellent energy efficiency on participants’ mobile devices over the 24 to 72 hour study periods. As shown in Table 2, the battery usage of Crepe ranged from 0%, 0.91% to 14%, with most participants reporting usage of 2% or less. This indicates that running Crepe in the background for data collection has minimal impact on the battery life of participants’ smartphones, making it a power-efficient tool for mobile data collection studies.
However, we observed a limitation of Crepe during the study: the background process management on Android can terminate Crepe to conserve resources. This is a known issue for similar mobile data collectors (Ferreira et al., 2015). For P1, we suspect the Android operating system terminated Crepe in the background, resulting in data loss at around hour 30. Only after we noticed it and prompted P1 to check the running status of Crepe did the data resumed to be collected. To address this issue, for later participants, we improved Crepe’s “heartbeat” mechanism and had the app send a heartbeat signal every 10 minutes. Once we noticed a potential inactive status of Crepe, we sent reminder emails to participants and asked them to open Crepe and check its running status. In future iterations of Crepe, this heartbeat mechanism could be combined with a dashboard to alert researchers when data collection becomes offline, enabling timely interventions to minimize data loss.
5. Discussion
5.1. Collection Support Tools for De-centralized Data Access
Crepe complements existing open-source data collection tools like AWARE (Ferreira et al., 2015) and Purple (Schueller et al., 2014) that focus on collecting sensor data, and others like ShiptCalculator (Calacci and Pentland, 2022) that enable workers to share screenshots of their pay history for wage transparency advocacy. Crepe targets the generalized collection of various types of mobile screen data. This contains great potential for more academic researchers to more easily access datasets and uncover more quantitative, analytical insights. Crepe aligns with the growing movement towards a more open and equitable ecosystem in digital economy (Arrieta-Ibarra et al., 2018).
5.2. Careful Considerations for User Privacy
The collection of user data through Crepe raises important privacy concerns regarding users’ app usage and screen information. Throughout the design and implementation of Crepe, user privacy has been a top priority, ensuring real-time transparency of collected data and employing high-standard encryption for data security. It is crucial to emphasize that Crepe is intended solely for research purposes, under scenarios where researchers obtain explicit consent from participants to collect screen data. In the future, we plan to continuously enhance the privacy-related features in Crepe. For instance, prior to the public release of Crepe, we intend to include a feature that hashes personal identifiers before sending them to researchers, thereby protecting users’ privacy, similar to the approach taken by previous data collectors (Ferreira et al., 2015). There have also been differential-privacy-based approaches that obfuscate sensitive data on collected mobile app UI snapshots (Li et al., 2020), which we can adopt in Crepe. We acknowledge that protecting users’ privacy in data collection remains an open research question, and we encourage the research community to collectively work towards improving our ethical standards in this regard.
5.3. Implications for Data Labor and Data Governance
Our approach for the decentralization of data access has important implications for data governance and the broader ecosystem of data-driven technologies. The rise of digital platforms has led to the concentration of user data in the hands of corporations and platforms, creating a “data monopoly” that limits the potential benefits of academic research and user-centric data-driven technollgies (Jia-Jun Li et al., 2022). By breaking the data monopoly and providing users with access to their own data, our approach can help bridge the data divide and empowers individuals and communities to leverage their own data for personal and collective benefits, rather than being solely exploited by corporations.
This increased transparency can help to identify and address biases and inequalities in the algorithms (Panch et al., 2021; Kordzadeh and Ghasemaghaei, 2022). By providing access to users’ own data on platforms, distributed, user-centric data-driven technologies can be developed in user groups and communities to shift the focus of benefits from corporations and platforms to end users. This aligns with the vision of a more equitable data ecosystem, where the value generated by user data is distributed more fairly among stakeholders (Posner and Weyl, 2018).
6. Limitations and Future Work
While Crepe offers a novel approach to mobile screen data collection, we acknowledge it still contains a few limitations. One issue is that Crepe can be terminated by the system in the background after a period of inactivity, potentially leading to data loss. This is a known issue for existing mobile data collectors (Ferreira et al., 2015). Additionally, system or app updates can potentially change the UI hierarchy of app screens, invalidating pre-defined graph queries and requiring researchers to update them during the study to ensure compatibility. Running an Android application continuously in the background without exhibiting malware-like behavior presents a challenge. Yet, we plan to refine Crepe’s notification system and collector active status to account for scenarios where Crepe gets killed in the background processes by the Android system. In addition, a few directions can also be explored to expand the capabilities of Crepe. Firstly, developing visualization dashboards for both researchers and participants would enable real-time monitoring of the collection status and the data contributed. Secondly, empowering users with the ability to delete all collected data after dropping out of a study would enhance their control over their data contributions and agency.
Currently, Crepe’s focus is on screen textual data, with great potentials in incorporating more rich multimodal data, such as screenshots and user interactions. However, we also acknowledge that there are certain limits to the types of data our approach cannot collect. For example, information rendered in game or 3D engines are hard to collect, since they do not show up in Android Accessibility Services’ responses. In addition, the expressiveness of our designed Graph Query also has boundaries. In certain screen structures, Graph Query may default to collecting data based on their absolute screen location, limiting the flexibility.
We plan to open source Crepe to promote the ethical use of our novel approach in future academic research data collection. We also welcome contributions from the research community to Crepe to make it more ethical, robust, and adaptive to various data collection scenarios.
7. Conclusion
In this paper, we introduce Crepe, a novel mobile data collection tool that enables researchers to easily define and collect user interface screen data from participants’ Android devices. Crepe utilizes Graph Query, which allows flexible identification, location, and extraction of target UI elements. The tool’s key feature is its low-code, programming by demonstration approach that lets researchers specify data to collect through simple interactions with example app screens. A user study with 7 participants demonstrated Crepe’s ability to effectively collect screen UI data across different devices, while identifying areas for future improvements. Overall, Crepe represents an important step towards democratizing mobile app data for research by reducing technical barriers and giving participants more agency. Code for Crepe will be open-sourced to support the academic community in future research data collection.
Appendix A Appendix
A.1. Details of Our Best Query Selection Prompt
As described in Section 3.2.2, we used large language models to select the best Graph Query that most accurately matches the researcher’s data collection intention. Specifically, our prompt uses the following structure:
We defined this new graph query to locate relevant data on mobile UI structure. Now, we have generated the following potential queries:
  1. (1)
    …(first query)
  2. (2)
    …(second query)
  3. (3)
    …(etc.)
Which one matches the following description the best?
[User description of the target data]
Please only return the index number, without any additional information, even periods or comma.
Example response format: 0
Example generated graph query candidates can be found in Figure 3 section 2. In the user description of the data to collect, in our user interfaces, we encourage the user to explain the target data’s relations with other content on screen, e.g., “Instagram Story advertisement title, the brand name above text Sponsored”.