### v2

# Crepe: A Mobile Screen Data Collector Using Graph Query

Abstract.
Collecting mobile screen information datasets remains challenging for academic researchers. Commercial organizations often have exclusive access to mobile data, leading to a “data monopoly” that restricts academic research and user transparency. Existing open-source mobile data collection frameworks primarily focus on mobile sensing data rather than screen content. We present Crepe, a no-code Android app that enables researchers to collect information displayed on screen through simple demonstrations of target data. Crepe utilizes a novel Graph Query technique, which augments mobile UI structures to support flexible identification, location, and collection of specific data pieces. The tool emphasizes participants’ privacy and agency by providing full transparency over collected data and allowing easy opt-out. We designed and built Crepe for research purposes only and in scenarios where researchers obtain explicit consent from participants. Code for Crepe will be open-sourced to support future academic research data collection.
journalyear: 2024
1. Introduction
To conduct research on mobile information consumption and behavior, researchers often turn to screen data, the rich information displayed on users’ mobile interfaces, as a window into the information that algorithms show people and how people interact with technology in everyday lives. Researchers analyze mobile device screen info datasets to uncover insights in Human Computer Interaction (HCI), social sciences, and adjacent fields (Raento et al., 2009). Unlike traditional mobile sensing data that captures device states and sensor readings, screen information represents what users actually see and engage with, making it an invaluable resource for understanding the user side of interactions (Brown et al., 2013; Rogers and Marshall, 2017; Bentley et al., 2016).
However, collecting mobile screen data has remained frustratingly elusive for academic researchers. In practice, access to comprehensive mobile app usage data is often exclusively controlled by smartphone platforms like iOS and Android, or by individual app developers who guard their user data closely (Chandramouli et al., 2015; Generosi et al., 2020). Academic researchers find themselves in a challenging position: they must either rely on limited public developer APIs—which are subject to corporate policy changes and restrictions (Reddit, 2023; Barnes, 2023)—or negotiate complex collaborations with commercial organizations that may compromise research independence (Kramer et al., 2014). This “data monopoly” creates significant barriers to empirical mobile research (Jia-Jun Li et al., 2022).
However, many researchers have limited resources and technical expertise to build their own data collectors. Meanwhile, the availability of developer APIs are subject to corporations’ discretionary policies (Reddit, 2023; Barnes, 2023). When researchers collaborate with commercial organizations to access mobile data (Kramer et al., 2014; Deka et al., 2017), they often need to align with the organizations’ business goals and public images, restricting the independence of academic research (Kramer et al., 2014). These phenomena are recognized as commercial organizations’ “data monopoly” in previous research (Jia-Jun Li et al., 2022) and pose challenges for empirical studies of mobile screen data.
While numerous open-source mobile data collection frameworks exist to support academic research, most mature solutions focus on mobile sensing data rather than the actual screen information that users see and interact with (Ferreira et al., 2015; Aharony et al., 2011; Cardone et al., 2013). The technical challenge lies in reliably and automatically identifying which user interface screens contain target data, then locating and collecting that specific information. Current approaches often require users to actively input data or upload screenshots manually (Calacci and Pentland, 2022), or they continuously record user screens—an approach that is both inefficient and raises serious privacy concerns (Krieter, 2019).
In this work, we ask: how can we democratize mobile screen data collection for academic researchers, while respecting user privacy and maintaining data quality?
We created Crepe [^1] Crepe is an acronym for “Collector for Research Experiments of Participant Experiences”, a novel Android data collector app that embodies a programming-by-demonstration approach to mobile screen data collection. At its core, Crepe introduces Graph Query, a query language that augments mobile UI screen structures to support flexible identification, location, and collection of specific data pieces. With Crepe, researchers can define what data to collect through simple, no-code demonstrations, while participants maintain full transparency and control over their data through intuitive privacy controls and real-time collection feedback.
We designed Crepe with participant agency at its center. Users see exactly what data is being collected through semi-transparent overlays and dedicated data pages, and they can leave any study at any time by simply removing the collector from their app. This approach transforms mobile data collection from an opaque, researcher-controlled process into a transparent, participant-empowered collaboration.
To evaluate Crepe, we conducted a series of three user studies with both researchers and participants, uncovering performance characteristics, limitations, and future directions in both laboratory and real-world settings.
In all, our contributions include:
  1. (1)
    A novel query language, Graph Query, to reliably identify, locate, and collect target data on UI screens;
  2. (2)
    Crepe, the first Android data collector app to support flexible, customizable, low-code UI screen data collection using programming by demonstration (PBD) (Cypher and Halbert, 1993);
  3. (3)
    A series of three user studies demonstrating Crepe’s effectiveness in empowering both researchers and participants in mobile screen data collection.
2. Background and Related Work
2.1. Understanding Human Activities via Mobile Data
Mobile devices, especially smartphones, are now ubiquitous in everyday lives. Mobile applications harness various information about user social activities, which are embedded in users’ interactions, app usage patterns, and preferences. As a result, researchers in HCI, and social science have been leveraging rich data gained from mobile apps to better improve user experiences (Deka et al., 2017; Wang et al., 2023), uncover societal trends (Vieweg et al., 2010; Starbird et al., 2014; Starbird and Palen, 2010), and understanding user behaviors and attitudes (Fakhradyan, 2021; De et al., 2017; Balaban et al., 2022; Kramer et al., 2014).
Screen data from mobile phones, i.e. information displayed on mobile screens, are particularly valuable for understanding user behavior and preferences. Screen data capture visual information and layout of the apps, providing insights into how users navigate and engage with different features. By analyzing this rich, granular data, researchers can gain a deeper understanding of user needs, preferences, and pain points, which can inform the design of more user-friendly and engaging mobile experiences (Ballard, 2007).
2.1.1. Analyzing Behavioral Data Streams
Researchers from diverse backgrounds, such as social science or business, mine mobile data for behavioral patterns and insights. Experience Sampling Method (ESM) is a research methodology to obtain in situ data for constructing an understanding of subjects’ daily behaviors, feelings, and thoughts of participants (van Berkel et al., 2017; Yue et al., 2014). In the survey conducted by Berkel et al., smartphone-based ESM benefits researchers in improving data quality through validation, collecting rich multimodal data for context reconstruction, and enabling real-time data collection and analysis (van Berkel et al., 2017). ESM can provide social scientists with combined human and sensor data in addition to background logging (Raento et al., 2009). Aside from sensor data, ESM questionnaires can also capture external aspects of experience, such as time, place, activities, and companion (Hektner et al., 2007). For instance, Yue et al. adopted ESM to study the function of photos. They found that photos can be used to trigger memory during follow-up interviews and as a beneficial component in data analysis (Yue et al., 2014). In this work, our goal is to create a low-code mobile UI screen data collection tool. Our app Crepe can facilitate academic researchers to instrument screen data collections without the need to extensively develop customized data collectors, or have public access to data APIs.
2.2. Computational Understanding of Mobile UI Structures and Data
Technical researchers have developed intelligent systems that achieve computational understanding of mobile user interfaces. Sugilite, a mobile PBD system that the execution of various tasks through user-driven multimodal direct manipulations, parses UI elements and layout information to understand the intentions contained in user activity (Li et al., 2017). The Sugilite system leverages the accessibility API of Android to encapsulate high-quality relevant behavioral data on mobile devices. UI understanding and summarization is another topic that HCI researchers are interested in. An accurate and succinct description of UI semantics can facilitate seamless interactions as well as bridge the gap between language and user interface. Prior research such as Screen2Vec (Li et al., 2021) and Screen2Words (Wang et al., 2021), adopted Machine Learning (ML) techniques to support UI design. More recent approaches have also leveraged the advance of Large Language Models (LLMs) and large multimodal models to enhance conversational interactions on mobile UI (Wang et al., 2023; Yan et al., 2023). In this work, with our proposed novel Graph Query, we aim to create a deterministic and reliable query language that can reliably and deterministically identify and locate a piece of target information on screen. This compliments existing research on mobile UI understanding in addition to existing deep-learning-based mechanisms.
2.3. State-of-the-art Mobile Data Collection Tools
A variety of data collection tools and frameworks have been proposed to support academic data collection of various types of data. Aware is a mobile instrumental tool designed for collecting usage context through sensors on mobile devices (Ferreira et al., 2015). Rather than being a simple mobile data collector, Aware creates a collaborative framework for researchers to share context data with the community. Purple Robot is an extensible and modular development platform that supports behavioral and clinical intervention (Schueller et al., 2014). The tool enables Behavioral Intervention Technology (BIT) stakeholders to create mobile apps that collect user management, content authorship, and content delivery data, through which physician-scientists can evaluate and share data resources to increase knowledge finding and generalization. ShiptCalculator is a research tool designed to track and share workers’ aggregated data about their pay to empower workers in offering awareness of wage transparency and advocacy to campaigns. In the data collection process, Workers send a screenshots of their pay history to ShiptCalculator, which parses them using OCR and stores structured data. The system then sends a validation text to workers and allows them to explore their pay details  (Calacci and Pentland, 2022). However, no current tool supports the flexible collection of data displayed on mobile UIs based on our review of previous work. Our design of Crepe in this study aims to solve this exact problem.
3. Proposed Method
Our goal for the Crepe data collector is three-fold. First, we seek to empower researchers lacking programming expertise to easily create and deploy data collections. Second, we prioritize the privacy and agency of individual data collection participants. Lastly, we also want to address the following observed technical challenges in reliable mobile screen data collection, which hinder even technical researchers:
  1. (1)
    The difficulty in automatically detecting that target data showed up on screen;
  2. (2)
    The lack of precision in locating target data and accessing its content;
  3. (3)
    The limited generalizability for dynamically changing data content.
To solve these challenges, previous collectors have asked participants to either directly input target data or upload screenshots or screen recordings containing target data (Calacci and Pentland, 2022; Krieter, 2019). However, these practices significantly increase participants’ efforts and reduce the reliability of the collected data’s quality.
In contrast, in Crepe, we used a novel Graph Query technique to address the above challenges. Graph Query has the following characteristics:
  1. (1)
    Triggering a data collection only when the target data shows up;
  2. (2)
    Locating and accessing only the target data on screen;
  3. (3)
    Generalizing easily to diverse data types, especially dynamically changing data.
For the rest of this section, we first give an overview of the experience design of Crepe, our Android mobile data collection app. Then, we dive into the technical details of the novel Graph Query we designed and implemented as the backbone of Crepe.
3.1. The Crepe Data Collector User Experience
Figure 2. The workflow of Crepe for our main user groups: data collection researchers and participants. Researchers create a new data collector by demonstration A and share the collector ID with participants B. Participants add the collector to their own devices C, which runs in the background to collect the specified target data D. Note that in step D, Crepe uses Android Accessibility Service to access the view hierarchy of the current UI screen, then processes the view hierarchy, instead of directly working with screenshot images. The collected data is transmitted to a database E for the researcher to analyze and gain insights F. The colors in the background indicate each user group’s experience involved in the holistic Crepe pipeline.
[…]
The figure illustrates the workflow of the Crepe system with 6 labeled steps: A) The researcher creates a new data collector by specifying the time period, app package, data field and description through a graphical interface on their device. B) The researcher shares the collector ID with participants. C) Participants add the collector to their own devices by entering the shared collector ID into the Instagram app running on their phone. The collector details like start/end time are shown. D) On the participant’s device, the Crepe app automatically identifies and locates the target UI data elements (like “walmart”, “deucelove”) using Graph Query, based on the UI screenshot. E) The target data elements collected by Crepe on participant devices get transmitted to a database. F) The researcher analyzes the collected data from the database to gain insights for their study. Arrows show the flow of information between the researcher device, participant devices, and the data collection database.
The Crepe data collector has two main target user groups: researchers who want to instrument a mobile data collection study, and study participants who want to contribute mobile data to ongoing data collections [^2] We intend Crepe to be used for research purposes only where participants give their explicit consent to researchers before data collection begins..
Researchers have two main tasks when using Crepe (Figure 2): creating a collector A, and sharing collectors with study participants B. A researcher can use Crepe’s graphical user interfaces to define the start and end dates of a collector, as well as the target app to collect data from.
The most important yet trickiest step in defining a data collector is specifying the target data on screen to collect. In Crepe, our design goal is to minimize the complexity of this specification process for researchers, by only asking them to tap on the screen data to collect in the target app, and describe their intention to collect these data (Figure 2 A). Crepe automatically translates these two pieces of information into a formal, executable Graph Query, which includes information about the target data’s parent app package, component type, and relations with other UI entities on screen. This Graph Query is the core of a data collector and will be used to reliably identify and precisely locate target data on the study participants’ devices. This specification process follows a programming-by-demonstration paradigm (Cypher and Halbert, 1993) and provides a simple, no-code experience for researchers to specify the target data. (The technical detail will be further explained in the following subsection on Graph Query.) After the target data are specified and descriptions are added, a collector is created. Afterward, the researcher can copy and share the unique ID of this collector with participants B. The collector will run on the participants’ device for the specified data collection periods. Afterwards, the researcher can receive the data collected in a self-defined database (E) and analyze it to uncover research insights (F). Our initial implementation of Crepe transmits collected data to a pre-configured Google Firebase Realtime Database that encrypts data both in transit and at rest [^3] https://firebase.google.com/support/privacy. We plan to support researchers’ customized configuration of their own database servers in the future.
For study participants, they can easily add a collector to Crepe by entering the unique collector ID shared by the study’s researchers C, using Crepe’s graphical user interfaces. After granting Crepe with the necessary Android Accessibility Service permission, the added collector(s) will start automatically running in the background of the participants’ devices (D). Guided by the Graph Queries associated with the participants’ added collectors, Crepe will selectively seek the appearance of target data exclusively when the designated target apps are active. Once the target data is detected, Crepe will scrape and transmit the relevant information to the researchers (E). The participant will be able to monitor the active status of the collector(s) they have joined in Crepe. To drop out of a study, participants can simply delete a collector from their Crepe app, and the collector will no longer run in the background.
Figure 3. Two design features that enhance data collection transparency for system users (data contributors). When the system operates in the background, a yellow highlight appears over the collected content at each time of collection ( 1). Additionally, we designed and implemented a page that displays all collected data organized by collector, allowing users to review their contribution history and details ( 2).
[…]
This image displays two key transparency features of a data collection system interface: The left side shows multiple overlapping interface elements, including a map in the background, an Uber ride selection menu, a news article with a headline about cryptocurrency, and a Facebook promotional card. A yellow highlight surrounds the active content being collected, drawing attention to what data is being captured. This section is labeled “1 Visual highlight for collected data.” The right side shows a data dashboard interface titled ”My Data” that provides users with a detailed log of their collected data. The dashboard displays information for two data collectors: Instagram and Uber. For each collector, it shows when collection started (January 29, 2025 for Instagram, February 1, 2025 for Uber), the scheduled end date (February 8, 2025 for both), and the total number of data entries (46 for Instagram, 19 for Uber). The dashboard also displays individual data collection events with timestamps, event types, and content descriptions, including specific details about Uber rides such as fare information, pickup times, and driver proximity. This section is labeled “2 Data dashboard displaying all collected data.” Together, these features provide transparency by highlighting data as it’s collected and giving users access to a comprehensive record of their contributed data.
Data collection transparency
Ensuring a transparent communication of data collection is critical in our design of Crepe. We ensure participants have the full control over their participation status and data contributions using two main measures. First, we adopt an “opt-in” participation mechanism, where participants actively add collectors to their Crepe app to join data collections. We intend this to be combined with the explicit consents that researchers must obtain before each data collection study starts. Second, each participant can opt out of any collection, at any time, by deleting the collector from their Crepe app, which will immediately terminate the collection on their mobile devices.
We also prioritized data collection transparency in Crepe design Figure 3. Every time data is collected from the user’s screen, a semi-transparent yellow highlight will appear over the collected data every time ( 1). We also designed a dedicated screen to show the users all of the data collected ( 2). We will also continuously release new privacy-centric features, such as daily or weekly push notifications summarizing data contributions, and the ability to request all their data be deleted when they withdraw from a study.
3.2. Graph Query
3.2.1. Building blocks of Graph Queries: screen entities and their relations
Figure 4. The detailed process of Graph Query generation and execution in Crepe. Section A shows how to generate a Graph Query: Crepe first uses Android Accessibility Service to access the view hierarchy of a UI screen 1 and augment it into a UI Snapshot graph, using UI elements’ characteristics and their relations. To identify the target UI element within the UI Snapshot graph, Crepe combines a set of UI element characteristics on this UI Snapshot to uniquely identify the target UI element on this UI Snapshot. We put these characteristics together to construct our Graph Query 2 and rank them based on their flexibility. In Section B, step 3 depicts the execution of a chosen Graph Query on two new screens, one containing the target data and one not. The set of unique characteristics of the target UI element ensures we only successfully collect the target data in “duecelove”, but not on the other screen.
[…]
The figure consists of two main sections, A and B, presented side by side. Section A, labeled “Query Generation”, displays a UI snapshot from a mobile app with various GUI elements like buttons, images, and text. It shows the process of creating Graph Queries by combining the relations between these screen entities. Section B, titled “Query Execution”, demonstrates the execution of a selected Graph Query on different UI snapshots. The query successfully identifies and locates the target element “duecelove” on one snapshot but fails to find a match on another screen. The figure uses visual cues like arrows and graphs to illustrate the flow of query generation and execution.
We develop a novel Graph Query in Crepe for accurate and consistent extraction of various types of screen information (Figure 4). The intuition behind Graph Query is to uniquely identify and locate the element(s) that contain the target data using the UI elements’ own attributes and relations with other elements on screen. Graph Query builds on top of the XML mobile screen hierarchy, similar to the DOM tree of web HTML, that is accessible through Android Accessibility Service. This mobile screen hierarchy contains Views on the current screen (1), each of which contains rich properties including content, screen location, etc.
First, we enhance the default mobile screen hierarchy provided by Android Accessibility Service into an augmented UI Snapshot, which contains additional relations between screen entities (Figure 4, UI Snapshot section). We defined a screen entity to be either an UI element (i.e., Android View) or an attribute of the UI element, in data representations such as strings or integers. The relations between these screen entities include:
  1. (1)
    UI elements to their own Android implementation attributes:
    • e.g. hasClassName, hasScreenLocation, isEditable
  2. (2)
    UI elements to their semantic content:
    • e.g. hasText, containsEmailAddress, containsMoney, containsDate
  3. (3)
    hierarchical relations between UI elements in the XML screen hierarchy:
    • e.g. hasParent, hasChild, hasSibling
  4. (4)
    spatial relations between UI elements on screenshot images:
    • e.g. right, left, above, below, near
These relations comprehensively encapsulate the various relations and semantic contents tha are useful for identifying and locating a date on screen. With such additional information, an UI Snapshot is a collection of subject-predicate-object triples denoted as (s, o, p), where s and o are two entities and the p is the relation between s and o.
3.2.2. Automatic generation of Graph Queries
When a researcher demonstrate the target data to collect by tapping on the it on screen (e.g. tapping on “apple” on the screen in Figure 4 Step 1), we locate the associated UI element in the UI Snapshot, and utilize a unique set of its characteristics in the UI Snapshot to generate Graph Query candidates (Figure 4, Step 2).
The Graph Query can be connected via 3 logical operators: conj (and), or (or), and prev (previous). We defined a set of context-free grammars, summarized in Table  1, to construct formal Graph Queries based on screen entity relations in UI Snapshots.
Expression Rule
E $\rightarrow$ e;
E $\rightarrow$ S;
S $\rightarrow$ (join r E);
S $\rightarrow$ (and S S);
T $\rightarrow$ (ARG_MAX r S);
T $\rightarrow$ (ARG_MIN r S);
Q $\rightarrow$ S — T;
Table 1. Context-free grammars (CFGs) for constructing Graph Query in Crepe. Q denotes the initial non-terminal symbol, terminal e denotes a GUI object entity, and terminal r denotes a relation. Other non-terminal symbols are employed for intermediary steps in the derivation process. We used this set of CFGs to automatically construct formal Graph Queries from the screen entity relations in UI Snapshots.
[…]
This figure presents a table outlining the context-free grammar (CFG) rules used to construct Graph Queries in the Crepe system. The table has two columns: the left column lists grammar expressions, and the right column shows their corresponding production rules.
  • E can be replaced by either ‎‘e;‘ or ‎‘S;‘
  • S can be replaced by ‎‘(join r E)‘ or ‎‘(and S S)‘
  • T can be replaced by ‎‘(ARG_MAX r S)‘ or ‎‘(ARG_MIN r S)‘
  • Q, the initial non-terminal symbol, can be replaced by either ‎‘S‘ or ‎‘T‘
Specifically, Graph Queries encapsulate combinations of the screen entity relations for the target UI element and Crepe creates combinations can uniquely identify only the target UI element on screen (Figure 4, Step 2). The combinations of screen entities are empirically set through experimentation. All Graph Queries will have hasPackageName and hasClassName to establish the context. In most cases, more than one Graph Query can uniquely locate the target UI element on the screen.
3.2.3. User Intent Disambiguation: Selecting the Best Graph Query
As described above, for each target data, Crepe generates a few Graph Query candidates that can each uniquely match the target data. However, some candidates are more generalizable to other screens—for example, to collect Instagram story ads, a query that collects the text above “Sponsored” is more likely to collect all target data, than a query that collects text that shows up at a specific screen location (the first and third queries in Figure 4, Step 2). We created a set of heuristics to rank the generalizability of generated queries; however, through our user studies, we learned that it is best to engage the researcher in this selection process.
Through iterative design with researchers (details in Section 4.2) and taking inspirations from past research (Li et al., 2018), our final design presents the top Graph Query candidates in natural language phrases for the user to select from. We translate Graph Queries into natural language phrases, such as “the UI element above the Button that says Sponsored”, to show to the user instead of its original query language form (Figure 4, Section 2, option 1). We translated the Graph Queries so that the researchers, especially the non-technical ones, are not exposed to unnecessary technical details of Crepe. We translate the Graph Queries using a large language model [^4] We used GPT-3.5-turbo in our implementation due to its ease of access, short inference time, and empirically high-quality outputs observed on our task. Our prompt can be found in Appendix A.1.. This ensures the selected Graph Query best reflects the researcher’s data collection goals.
3.2.4. Graph Query execution on UI Snapshots: retrieving the target data
After a Graph Query is created, it can be used to collect target data on any UI Snapshot with elements sharing the same set of entity relations. In the background, Crepe runs the Graph Query every time a relevant UI snapshot is updated by changes in screen content (Figure 4, Section B). We also limit the Graph Query to only run in the target app package, in order to effectively reduce unnecessary battery usage of Crepe. After identifying and locating the target UI element on screen (Step 3), Crepe can collect the target data to collect through the element’s attributes such as text and content description.
3.2.5. Summary
Graph Query showcases three major strengths in screen data collection. First, the data collection mechanism using Graph Query is fully deterministic and thus much more reliable than alternative solutions like using Optical Character Recognition (OCR) or deep-learning based detection models. Second, Graph Query provides much more flexibility in identifying and locating target UI elements through their implementation attributes, semantic content, hierarchical screen structure position, and screen location relations (Section 3.2.1). For instance, a researcher can utilize the pre-defined hasPrice relation to collect Uber drivers’ task payment information, a common task on gig worker support platforms (Calacci and Pentland, 2022). In the future, developers can also extend existing entity relation types for more customized data collection tasks. Lastly, by utilizing Android’s Accessibility Service, using Graph Query can easily integrate the collection of user interaction data, a feature included in Crepe yet unsupported in most vision-based data collection tools. In the future, Crepe can be extended to collect more rich media information such as screenshots under the study participants’ consent555 Link to the documentation of Android Accessibility Service’s API for taking screenshots..
3.3. Implementation
Crepe was implemented using Android, Google Firebase Realtime Database, OAuth, and OpenAI API. Note that we only used OpenAI large language model for Graph Query translation (Section 3.2.3) and the whole data collection process does not involve any deep learning. The app was developed in Java using Android Studio and is compatible with mobile devices running Android 9.0 or above (around 95.4% of the Android mobile phone market as of February 2025666https://composables.com/android-distribution-chart). Firebase Realtime Database was selected due to its high-standard data encryption both in transit and at rest, real-time data synchronization, and reliable user authentication. To implement the query language, Crepe leverages the Accessibility Service and captures Accessibility Events as a result of changes in the content of the screen. The captured Accessibility Event contains the new UI screen hierarchy [^7] https://developer.android.com/reference/android/view/accessibility/AccessibilityEvent. The code for Crepe will be open-sourced to enable easy adoption and further development within the research community.
4. Evaluations
To test and validate Crepe, we conducted a series of three evaluation studies. The first two studies focused on researchers’ experience to collect data (Study 1) and Crepe’s data collection accuracy (Study 2). The third study investigated Crepe’s battery usage and perceived interruptions in the wild for participants (i.e. data contributors) (Study 3). The first two studies were designed to be in-lab, while Study 3 was conducted as a field study. These three studies together aim to answer the following research questions:
  1. (1)
    RQ1: How well does Crepe support researchers in collecting mobile screen data (Study 1)?
  2. (2)
    RQ2: How accurately does Crepe collect target data on participants’ screens (Study 2)?
  3. (3)
    RQ3: How does Crepe influence device performance for participants (i.e. data contributors) in a real-world data collection scenario (Study 3)?
4.1. Evaluation Scenarios
We wanted our performance evaluation of Crepe to cover a range of different data collection needs. Looking broadly into previous research in HCI, CSCW, and Data Science, we extracted three scenarios where researchers can use Crepe to collect user screen data, Our evaluation studies were based in these scenarios.
Figure 5. We identified three potential usage scenarios of Crepe for our evaluation. Scenario 1 collects the advertiser information in Instagram Story ads, scenario 2 collects Uber ride pricing when user requests a ride, and scenario 3 collects the top feed item in Chrome Discover Feed recommendations. Crepe shows a yellow highlight overlay every time the target data is collected, providing full transparency to the collection participants.
[…]
This figure shows three mobile phone screenshots arranged horizontally, each labeled as a different scenario. In Scenario 1 (left), an Instagram Story advertisement is shown with the text “Inspiration is everywhere. Find yours.” along with several colorful images featuring someone in casual settings. In Scenario 2 (middle), an Uber app interface displays ride options near a Walmart Supercenter. A yellow highlight box surrounds ”UberXL” pricing information showing ”7 min away” and ”Affordable rides for groups up to 6.” In Scenario 3 (right), a Google Chrome mobile browser interface shows the Discover feed. A yellow highlight box surrounds a news article preview titled “This Is A Big Deal - Bitcoin And Crypto Now Braced For A Huge U.S. Price Earthquake” from Forbes published 7 hours ago. Each screenshot represents a different type of digital data that could be collected for research purposes, with the yellow highlighting indicating the specific elements of interest.
Scenario 1: Instagram Story Advertisements
Instagram Story advertisements are sponsored short image/video segments that appear between an Instagram user’s following content. They usually come with a “Sponsored” label to be distinguished from users’ regular following content. Mutiple research studies have been conducted in recent years around Instagram Story advertisements, to understand how users perceive and understand this specific form of advertising (Balaban et al., 2022; Haldborg Jørgensen et al., 2023; Fakhradyan, 2021). However, it is not easy to collect a dataset of Instagram story advertisements.
Getting access to a dataset of Instagram’s Story ads can enable research contributions in fields including social media research (Hanbazazh and Reeve, 2021), algorithm auditing (Ulbricht and Yeung, 2022; Eslami et al., 2015, 2016), and usable privacy (Chen et al., 2024). Specifically, we can empirically observe users’ interactions with Instagram Story ads, as well as inferring dwelling time over Story ads based on the first and last content change while a specific ad appears on screen. More generally, some broader questions we can answer with the dataset include: What empirical understanding of Instagram’s personalized advertisement algorithms can we obtain from quantitative data analysis? How do Instagram Story advertisements reveal the platform’s portraits of individual users’ preferences? And what are users’ folk theories of Instagram’s personalized ad algorithm?
Scenario 2: Uber Pricing
Ride-sharing platforms like Uber implement dynamic pricing strategies that adjust fare prices based on various factors including demand, time of day, and location. Collecting real-time price data for trip requests would allow researchers to study pricing algorithms, examine fare consistency across different user groups, and analyze how external factors influence ride-sharing costs. Previous research has attempted to understand these pricing mechanisms through qualitative measure (Lenjo, 2017) and quantitative analysis (Chao, 2019). However, current data sources cannot cover many user-specific aspects such as price variations across user groups and discount information. This data could help researchers investigate questions about algorithmic fairness in transportation access, analyze the relationship between surge pricing and local events, and study how pricing strategies affect user behavior in different geographical and temporal contexts. In our scenario, we focus on collecting the pricing information for participants’ Uber ride requests.
Scenario 3: Chrome Browser Discover Feed
Mobile browsers like Google Chrome integrate algorithmic news feeds into their search interface. These algorithmic feeds, often supported by deep learning, curates content based on users’ browsing behavior and inferred interests (Zheng et al., 2018). These algorithmic feeds, while work at an individual user level, have wide and profound social implications (Sandvig et al., 2014; Ribeiro et al., 2020).
Using Chrome as an example, accessing its Discover feed data from the user’s perspective enables research in algorithm auditing, helping to analyze content recommendation patterns, assess personalization mechanisms, and identify potential biases (Sandvig et al., 2014; Guess et al., 2023). Researchers have proposed audit studies to uncover discriminatory practices (Sandvig et al., 2014). Researchers can investigate how these algorithms shape information exposure by examining content diversity, frequency, and source representation. Understanding these dynamics in the Discover feed can identify biases impacting users’ information access and inform design improvements for more equitable content curation.
4.2. Study 1: Evaluating Crepe Usability with Researchers (In-Lab)
Study 1 evaluates how Crepe effectively supports researchers. While Crepe serves both researchers (collector creators) and users (data contributors), this study focuses specifically on the researcher experience during collector creation. We examine the usability of our Programming By Demonstration (PBD) approach in creating a data collector and try to understand what are researchers’ potential use cases of Crepe.
4.2.1. Procedure
Each one-hour session began with a brief introduction to Crepe’s motivation and core features. Participants then created data collectors for three representative usage scenarios (detailed in Section 4.1) using Crepe’s interface. After completing the tasks, participants completed a modified System Usability Scale (SUS) questionnaire (Brooke, 1996), including five-point Likert scale questions evaluating Crepe on its usability and utility for future research projects. We concluded with semi-structured interviews exploring participants’ envisioned uses of Crepe in their research. All studies were carried out in person and participants received a \$15 digital gift card for their time. To evaluate the accuracy of the PBD-generated Graph Queries, two authors independently analyzed whether each Query correctly captured the researcher’s specified data collection targets.
4.2.2. Participants
We recruited 5 HCI researchers through social media advertisements and word of mouth. All participants had experience with data collection and had interests in collecting users’ screen data for research purposes. Most participants have experience collecting various types of user data in the past (see participant demographics in Section A.2).
Figure 6. The usability questionnaire results for Study 1, evaluating researchers’ perception of Crepe’s usability and data collection capabilities. The tool received positive feedback, demonstrating its usability and utility for potential data collections studies.
[…]
A horizontal bar chart showing responses to 9 statements about Crepe’s usability on a 5-point Likert scale from ’Strongly agree’ to ’Strongly disagree’. Most responses trend positive, with ’Various functions are well integrated’ receiving the strongest agreement (5 respondents strongly agreeing). The chart uses a color scheme of dark blue for ’Strongly agree’ through pink for ’Strongly disagree’, with statement text aligned left and response counts displayed in bars.
4.2.3. Results
The usability evaluation showed positive results (Figure 6), with participants rating Crepe favorably on all questions. Most participants also showed a strong interest in using Crepe for future screen data collection studies and think it would not require much adaptation to Crepe.
In Study 1, as researchers created data collectors using Crepe, we also evaluated the programming by demonstration (PBD) pipeline for authoring candidate Graph Queries. Our evaluation was carried out on in total 15 created collectors (3 example scenarios for each of the 5 participants). We found that in all of the 15 occasions (15/15, 100%), our PBD pipeline successfully generated Graph Queries that correctly match users’ data collection goal (see Figure 4). Based on participant feedback, we iteratively improved the design of the user disambiguation process. We landed on our final design as described in Section 3.2.3, making it intuitive to pick the best Graph Query that matches researchers’ data collection goals.
Our participants identified several novel research opportunities enabled by Crepe’s screen data collection capabilities. P1 mentioned the interest in collecting Twitter (X)’s feed data to understand the platform’s feed algorithm. P2 would like to collect data from Uber to understand how ride prices change throughout a day and throughout different days of a week. P5 proposed to collect journal apps’ suggestion prompts to gauge what the app understands about users’ everyday activities. Interestingly, P1 also brought up the need to collect data based on certain conditions—for example, only collect social media feed under the “For you” tab instead of “Following”—which can further expand the utility of Crepe.
4.3. Study 2: Evaluating Graph Query Data Collection Accuracy (In-Lab)
Study 2 evaluates how accurately Crepe collects screen data in realistic app usage scenarios. While Study 1 focused on researchers’ experience creating data collectors, this study examines the technical performance of Graph Query in locating and extracting target data.
The nature of the study required us to conduct it as an in-lab study. Measuring the accuracy of Graph Query in data collection requires us to compare (1) the target data that showed up on users’ screens, and (2) the data Crepe is able to collect in the process. To obtain (1), we recorded participants’ screen during the in-lab study, to establish the baseline.
Metric Overall Instagram Uber Chrome
Recall 95.80% (506/528) 94.7% (269/284) 96.6% (115/119) 97.6% (122/125)
Precision 96.00% (504/525) 95.0% (268/282) 100% (115/115) 94.5% (121/128)
F1 Score 96.0% 94.8% 98.3% 96.0%
Table 2. The evaluation results for Crepe in the three app scenarios. Our in-lab study 2 demonstrates that Crepe performs well on on the three selected data collection scenarios. We discuss the observed edge cases that Crepe cannot handle well in Section 4.3.3.
[…]
This table presents performance metrics for a classification system across three digital platforms (Instagram, Uber, and Chrome) along with overall performance measures, where the total number of entries analyzed varies by platform with 528 entries overall, 284 entries for Instagram, 119 entries for Uber, and 125 entries for Chrome. The recall (true positive rate) measurements show 506 out of 528 (95.80%) overall, 269 out of 284 (94.7%) for Instagram, 115 out of 119 (96.6%) for Uber, and 122 out of 125 (97.6%) for Chrome. The precision (positive predictive value) results indicate 504 out of 525 (96.00%) overall, 268 out of 282 (95.0%) for Instagram, 115 out of 115 (100%) for Uber, and 121 out of 128 (94.5%) for Chrome. The F1 scores (harmonic mean of precision and recall) are 96.0% overall, 94.8% for Instagram, 98.3% for Uber, and 96.0% for Chrome, with Uber achieving perfect precision but slightly lower recall, while Chrome demonstrated the highest recall but the lowest precision, and Instagram maintained consistent performance across all metrics around 95%, resulting in strong overall performance across all platforms with F1 scores above 94%.
4.3.1. Procedure
We provided participants with a development phone with Crepe pre-installed and configured to collect data from apps in the three defined scenarios (Section 4.1). Each participant spent approximately one hour using these apps, allocating around 20 minutes per app. Participants were instructed to use each app naturally as they would in their daily life, while also feeling free to switch between apps when needed to mirror realistic usage patterns. All studies were carried out in person and participants received \$15 digital gift cards for each hour they participated.
To establish ground truth for evaluating collection accuracy, we simultaneously recorded the phone screen during the study sessions. An author labeled the screen recordings to identify all instances where target data appeared as the ground truth, and another author separately verified the results. We compared this ground truth data against the data collected by Crepe to compute collection accuracy.
4.3.2. Participants
We recruited 5 participants through social media advertisements and word-of-mouth. This provided approximately 240 minutes of naturalistic app usage data for our accuracy evaluation, generating a diverse set of UI states and interactions across different apps. Participant demographics and their app usage experience are detailed in Section A.3.
4.3.3. Results
Across all sessions, we recorded 528 data points across three apps as the benchmark (284 from Instagram, 119 from Uber, and 125 from Chrome). Crepe achieved an overall collection accuracy F1 score of 96.0% (precision: 96.0%, recall: 95.8%). Breaking down by app, collection accuracy F-1 scores were consistently high: Uber (98.3%), Instagram (94.8%), Chrome (96.0%). The minor variations in accuracy across apps can be attributed to differences in UI update frequencies and layout complexities.
Edge cases.
In the evaluation, we observed several scenarios where Crepe can make mistakes in data collection. First, some apps have UI layout variations for the same target data, requiring Crepe to set up multiple Graph Queries to collect the same data. In our evaluation, Instagram story’s ads are sometimes presented above a TextView containing “Sponsored”, and sometimes above a Button containing “Sponsored”. However, by demonstrating in both of these scenarios, researchers can still use Crepe to robustly handle such data collection tasks. In fact, we demonstrated three different UI layouts to Crepe during our evaluation study for Instagram Stories. Second, complex UI screens containing more than 500 UI elements can pose challenges to Crepe’s processing speed, potentially omitting data due to performance bottlenecks. We set a loose threshold to Crepe’s processing frequency to avoid repeatedly collecting the same data. We discuss some of our future improvement plans to address these challenges in Section 6.
In general, the results show that Graph Query can reliably collect screen data across various apps and UI screens. The collected dataset captured diverse scenarios including real-time price updates in Uber, dynamically positioned sponsored content in Instagram, and recommendation feeds in Chrome homepage. These results validate that Graph Query’s data collection mechanism is robust enough for real-world deployment while maintaining high accuracy across different apps and scenarios.
4.4. Study 3: Evaluating System Impact for Users (Field Study)
Now that we understand Crepe’s performance in lab, how well does it work in the wild? In Study 3, we conducted a field study with users who contributes their screen data through Crepe. Study 3 focuses on the real-world performance of Crepe, especially on its device performance and impact on data contributors’ regular phone usage experience. We designed Study 3 to evaluate how Crepe performs on different Android devices and OS versions, helping us catch any real-world deployment issues we couldn’t spot in the lab. Study 3 was conducted in example scenario 1 (Instagram story ads), since our target metrics will not significantly change when deployed for different apps’ data collections.
Participant Mobile OS Device Period Battery Usage
PC1 Android 14 Pixel 8 (own) 72 hours 1%
PC2 Android 14 Pixel 6 (provided) 72 hours 14%
PC3 Android 14 Pixel 6 (provided) 72 hours 9%
PC4 Android 13, OxygenOS 13.1 OnePlus 9R (own) 24 hours 0.91%
PC5 Android 12, Oxygen OS 12.1 OnePlus Nord (own) 24 hours N/A
PC6 Android 14 Pixel 6 (provided) 24 hours 2%
PC7 Android 9 Galaxy S10 (own) 24 hours 0%
Table 3. Mobile device usage data for study participants over 24 to 72 hour periods, including mobile OS, device model, and battery usage percentage. Three of the participants did not have an Android phone and used our provided device, with which they were asked to use Instagram following their regular habits during the study period. The participants decided the time period of the study based on their time schedule. Note that P5 reported that they could not found the battery usage item of Crepe on their phone Settings page after the collection. We were not able to help them find it either, despite our best efforts.
[…]
This table presents mobile device usage data for seven participants over study periods ranging from 24 to 72 hours. The table includes each participant’s mobile OS, device model, study duration, and the battery usage attributed to running the Crepe app. Three participants (PC2, PC3, and PC6) used provided devices (Pixel 6 with Android 14) because they did not own an Android phone. The remaining participants used their own devices, which varied in model and Android version. Battery usage varied by participant and device. PC1 used a Pixel 8 for 72 hours and reported 1% battery usage. PC2 and PC3, both using provided Pixel 6 devices for 72 hours, reported 14% and 9% usage, respectively. PC4 used a OnePlus 9R (Android 13, OxygenOS 13.1) for 24 hours and reported 0.91% battery usage. PC5, using a OnePlus Nord (Android 12, OxygenOS 12.1) for 24 hours, was unable to locate battery usage data in their device settings. PC6 used a provided Pixel 6 for 24 hours and reported 2% usage. PC7 used a Galaxy S10 (Android 9) for 24 hours and reported 0% battery usage. The results suggest that Crepe consumes minimal battery across a range of Android devices and OS versions. Most participants reported battery usage under 2%, indicating that Crepe is energy-efficient for background data collection. The only exception was PC2, who experienced 14% usage over 72 hours. A note below the table explains that PC5 could not find the battery usage stats despite assistance from the research team.
Participants
We recruited seven participants through social media advertisements and word-of-mouth. The participants are regular users of Instagram and Instagram Story. Each participant decided to contribute their data from 24 hours up to 72 hours based on their preferences and time schedule. We aimed at testing Crepe on various Android versions and device platforms, while also evaluating Crepe’s performance over short and longer collection duration. Four participants used their own Android devices. Since our main goal is not to collect a dataset that can be directly used in rigorous quantitative analysis but to empirically test our collector system, we also recruited three participants who are interested but do not have an Android device (P2, P3, P6). We provided each of them with an Android phone, with which they were asked to use Instagram during the period of the study, roughly following their regular Instagram usage habits. These three participants also contributed to our study by helping us assess the performance and usability of Crepe across different usage behaviors.
Study procedure
The user study followed a three-stage procedure [^8] The study protocol was reviewed and approved by the Institutional Review Board (IRB) at our institution.:
  1. (1)
    A 30-minute introductory interview to familiarize participants with the study and set up Crepe on their devices;
  2. (2)
    A testing period lasting between 24 to 72 hours, based on the participants’ time schedule and preferences, during which they were encouraged to use Instagram as usual.
  3. (3)
    A 30-minute concluding interview to administer a post-study questionnaire, collect feedback on participants’ experiences, and address any questions or concerns.
Throughout the study, we maintained email communications with participants to ensure Crepe functioned normally and conducted troubleshooting sessions when necessary. Each participant was compensated with 25 USD digital gift card for the data they contributed, and 20 USD for each hour of meeting sessions they attended. The participants’ participation duration, device information, and Crepe’s battery usage result after the study ended are shown in Table 3.
Figure 7. Temporal patterns of Instagram story ads collected in Study 3 using Crepe. This scatter plot displays unique data from participants P1–P3, who chose a 72-hour study period based on their availability. Circle size corresponds to the number of entries collected at each time point (shown by the number within each circle). The distribution reveals how ad exposure varied across participants and throughout the three-day period, with some participants experiencing more consistent ad delivery while others encountered more sporadic patterns. All data was passively collected while participants engaged in their regular smartphone usage. The 24-hour study results for the rest of participants is shown in Figure 8.
[…]
This image shows a scatter plot titled “Scatter Plot of Instagram Story Ads Collection (72-hour study)” with the subtitle “Dot size proportional to number of entries.” The plot displays data for three participants (P1, P2, and P3) over a 72-hour period. The y-axis represents the participants, labeled from P1 at the bottom to P3 at the top. The x-axis shows time in hours, ranging from 0 to approximately 70 hours. Each data point is represented by a colored circle where the size of the circle corresponds to the number of Instagram story ads collected, with larger circles indicating more entries. P1 (shown in blue circles) has data points scattered at hours 0 (8 entries), around hour 25 (1, 3, and 5 entries), and at hour 70 (1 and 2 entries). P2 (shown in purple circles) has more frequent data collection with clusters at hours 0-5 (with entries of 6, 2, 19, and 2), hours 10-15 (with 10 entries multiple times), hours 20-30 (with 4, 7, and 4 entries), hour 45 (with 10 entries), and hour 65 (with 2 entries twice). P3 (shown in green circles) has data points at hours 0-5 (15 entries), hours 10-15 (9 and 10 entries), hours 40-45 (2 entries twice), hours 55-60 (7, 1, and 1 entries), and hour 65 (7 entries). The legend in the top right identifies each participant with their corresponding color: P1 (blue), P2 (purple), and P3 (green).
Figure 8. The distribution of Instagram story ads collected in Study 3 using Crepe. This scatter plot represents unique data collected from participants P4-P7, who chose a 24-hour study period based on their availability. Each circle represents a collection event, with size proportional to the number of ads captured (indicated by the number inside each circle). Participants installed Crepe on their mobile phones and continued their normal daily activities while the system collected Instagram story ads in the background.
[…]
This image displays a scatter plot titled “Scatter Plot of Instagram Story Ads Collection (24-hour study)” with the subtitle “Dot size proportional to number of entries.” The plot tracks data collected from four participants (P4, P5, P6, and P7) over a 24-hour period. The horizontal axis represents time in hours (from 0 to 25), and the vertical axis shows the four participants. Each data point appears as a colored circle — olive for P4, orange-brown for P5, dark blue for P6, and light green for P7. The size of each circle corresponds to the number of Instagram story ads collected at that time point, with larger circles representing more entries. The specific number of entries is displayed within each circle. Participant P7 (light green circles) has the largest single collection of 35 entries at hour 0 and a smaller collection of 2 entries around hour 15. Participant P6 (dark blue circles) shows consistent activity throughout the 24-hour period, with multiple collection points at hours 0-2, 15-16, 18, 20, and 23. Participant P5 (orange-brown circles) shows activity primarily in the first 5 hours, with entries of 3-10 ads collected. Participant P4 (olive circles) shows a pattern of data collection at the beginning (hour 0), middle (hour 10), and end (hours 18-22) of the study period, with the largest collections of 24 and 28 entries occurring in the latter half of the study.
Study Results
After initial data analysis, Crepe was able to collect 358 Instagram Story advertisements from all seven participants, with an average of 51 ads per participant (Figure 7, Figure 8). The collected data reveals the number of unique Instagram Story Ads encountered by each participant at regular time intervals throughout their respective data collection periods. The data allow us to observe the frequency and distribution of Instagram Story Ads over time, potentially providing insights into the advertising strategies employed by businesses on the platform. We shared the data we collected from each participant with them individually upon request.
In the post-study questionnaire, which included two 5-scale questions, all participants unanimously reported that they felt minimal interruption and disturbance from Crepe while using their everyday apps and did not notice significant changes to their devices during the collection study. Here are details regarding these two questions and all participants’ response:
  • Q1: When using my everyday apps, I felt minimal interruption and disturbance from the Crepe app.
    • All participants responded 5 – Strongly agree
  • Q2: I noticed significant changes to my device during the data collection study.
    • All participants responded 1 – Strongly disagree
Based on the battery usage data collected in the post-study questionnaire, Crepe demonstrated excellent energy efficiency on participants’ mobile devices over the 24 to 72 hour study periods. As shown in Table 3, the battery usage of Crepe ranged from 0%, 0.91% to 14%, with most participants reporting usage of 2% or less. This indicates that running Crepe in the background for data collection has minimal impact on the battery life of participants’ smartphones, making it a power-efficient tool for mobile data collection studies.
However, we observed a limitation of Crepe during the study: the background process management on Android can terminate Crepe to conserve resources. This is a known issue for similar mobile data collectors (Ferreira et al., 2015). For P1, we suspect the Android operating system terminated Crepe in the background, resulting in data loss at around hour 30. Only after we noticed it and prompted P1 to check the running status of Crepe did the data resumed to be collected. To address this issue, for later participants, we improved Crepe’s “heartbeat” mechanism and had the app send a heartbeat signal every 10 minutes. Once we noticed a potential inactive status of Crepe, we sent reminder emails to participants and asked them to open Crepe and check its running status. In future iterations of Crepe, this heartbeat mechanism could be combined with a dashboard to alert researchers when data collection becomes offline, enabling timely interventions to minimize data loss.
4.5. Summary
Our three-part evaluation of Crepe examined its effectiveness in supporting researchers’ data collection needs (RQ1), accuracy in collecting target screen data (RQ2), and real-world performance impact (RQ3). Study 1 demonstrated that researchers could effectively use Crepe’s programming by demonstration approach, with all five participants successfully creating accurate data collectors for three different scenarios and rating the system’s usability favorably. Study 2 validated Crepe’s technical robustness through an in-lab evaluation, achieving high collection accuracy with an overall F1 score of 96.0% across Instagram, Uber, and Chrome applications, though some challenges emerged with complex UI screens containing over 500 elements. Finally, Study 3’s field deployment with seven participants showed that Crepe had minimal impact on device performance, with most participants reporting battery usage of 2% or less and unanimously indicating no noticeable interference with their regular phone usage. The system successfully collected 358 Instagram Story advertisements during the field study, though we identified opportunities for improving background process management to prevent occasional data collection interruptions. Together, these results demonstrate that Crepe effectively balances researcher needs, technical accuracy, and user experience in mobile screen data collection, while highlighting specific areas for future enhancement.
5. Discussion
5.1. Collection Support Tools for De-centralized Data Access
Crepe complements existing open-source data collection tools like AWARE (Ferreira et al., 2015) and Purple (Schueller et al., 2014) that focus on collecting sensor data, and others like ShiptCalculator (Calacci and Pentland, 2022) that enable workers to share screenshots of their pay history for wage transparency advocacy. Crepe targets the generalized collection of various types of mobile screen data. This contains great potential for more academic researchers to more easily access datasets and uncover more quantitative, analytical insights. Crepe aligns with the growing movement towards a more open and equitable ecosystem in digital economy (Arrieta-Ibarra et al., 2018).
The democratization of screen information access breaks the “data monopoly” that operating systems providers and app developers have long maintained (Jia-Jun Li et al., 2022). It shifts the data ownership back to end users, instead of leaving it only concentrated in the hands of tech platforms. By helping users see their own data, we can enable customized features that monitor and understand their behavior, supporting more informed device and app usage. This transparency allows individuals to make better decisions about their digital habits and identify patterns they might want to change.
By providing academic researchers access to screen data, they can better audit algorithms without depending on platform APIs that companies can (and often do) remove at will (Ribeiro et al., 2020; Sandvig et al., 2014; Calacci and Pentland, 2022; Kaplan and Sapiezynski, 2024). This independent access creates a more sustainable research ecosystem that isn’t vulnerable to corporate gatekeeping. Researchers can investigate questions about algorithmic bias, content recommendation systems, or user experience issues with greater freedom and thoroughness. This approach represents a fundamental shift in how we think about digital data ownership, treating screen data as something that belongs primarily to users rather than platforms. It encourages a more balanced relationship between tech companies, users, and the research community.
5.2. Careful Considerations for User Privacy
The collection of user data through Crepe raises important privacy concerns regarding users’ app usage and screen information. Throughout the design and implementation of Crepe, user privacy has been a top priority, ensuring real-time transparency of collected data and employing high-standard encryption for data security. It is crucial to emphasize that Crepe is intended solely for research purposes, under scenarios where researchers obtain explicit consent from participants to collect screen data. In the future, we plan to continuously enhance the privacy-related features in Crepe. For instance, prior to the public release of Crepe, we intend to include a feature that hashes personal identifiers before sending them to researchers, thereby protecting users’ privacy, similar to the approach taken by previous data collectors (Ferreira et al., 2015). There have also been differential-privacy-based approaches that obfuscate sensitive data on collected mobile app UI snapshots (Li et al., 2020), which we can adopt in Crepe. We acknowledge that protecting users’ privacy in data collection remains an open research question, and we encourage the research community to collectively work towards improving our ethical standards in this regard.
5.3. Implications for Data Labor and Data Governance
Our approach for the decentralization of data access has important implications for data governance and the broader ecosystem of data-driven technologies. The rise of digital platforms has led to the concentration of user data in the hands of corporations and platforms, creating a “data monopoly” that limits the potential benefits of academic research and user-centric data-driven technollgies (Jia-Jun Li et al., 2022). By breaking the data monopoly and providing users with access to their own data, our approach can help bridge the data divide and empowers individuals and communities to leverage their own data for personal and collective benefits, rather than being solely exploited by corporations.
This increased transparency can help to identify and address biases and inequalities in the algorithms (Panch et al., 2021; Kordzadeh and Ghasemaghaei, 2022). By providing access to users’ own data on platforms, distributed, user-centric data-driven technologies can be developed in user groups and communities to shift the focus of benefits from corporations and platforms to end users. This aligns with the vision of a more equitable data ecosystem, where the value generated by user data is distributed more fairly among stakeholders (Posner and Weyl, 2018).
6. Limitations and Future Work
While Crepe offers a novel approach to mobile screen data collection, we acknowledge it still contains a few limitations. One issue is that Crepe can be terminated by the system in the background after a period of inactivity, potentially leading to data loss. This is a known issue for existing mobile data collectors (Ferreira et al., 2015). Additionally, system or app updates can potentially change the UI hierarchy of app screens, invalidating pre-defined graph queries and requiring researchers to update them during the study to ensure compatibility. Running an Android application continuously in the background without exhibiting malware-like behavior presents a challenge. Yet, we plan to refine Crepe’s notification system and collector active status to account for scenarios where Crepe gets killed in the background processes by the Android system. In addition, a few directions can also be explored to expand the capabilities of Crepe. Firstly, developing visualization dashboards for both researchers and participants would enable real-time monitoring of the collection status and the data contributed. Secondly, empowering users with the ability to delete all collected data after dropping out of a study would enhance their control over their data contributions and agency.
Currently, Crepe’s focus is on screen textual data, with great potentials in incorporating more rich multimodal data, such as screenshots and user interactions. However, we also acknowledge that there are certain limits to the types of data our approach cannot collect. For example, information rendered in game or 3D engines are hard to collect, since they do not show up in Android Accessibility Services’ responses. In addition, the expressiveness of our designed Graph Query also has boundaries. In certain screen structures, Graph Query may default to collecting data based on their absolute screen location, limiting the flexibility.
Taking inspiration from previous PBD studies (Lieberman et al., 2000; Cypher, 1991), it is also helpful to ask users demonstrate on multiple target data examples. We will continuously incorporate these features into Crepe and welcome the research community to collaboratively improve it in the future. We plan to open source Crepe to promote the ethical use of our novel approach in future academic research data collection. We also welcome contributions from the research community to Crepe to make it more ethical, robust, and adaptive to various data collection scenarios.
7. Conclusion
In this paper, we introduce Crepe, a novel mobile data collection tool that enables researchers to easily define and collect user interface screen data from participants’ Android devices. Crepe utilizes Graph Query, which allows flexible identification, location, and extraction of target UI elements. The tool’s key feature is its low-code, programming by demonstration approach that lets researchers specify data to collect through simple interactions with example app screens. A user study with 7 participants demonstrated Crepe’s ability to effectively collect screen UI data across different devices, while identifying areas for future improvements. Overall, Crepe represents an important step towards democratizing mobile app data for research by reducing technical barriers and giving participants more agency. Code for Crepe will be open-sourced to support the academic community in future research data collection.
Appendix A Appendix
A.1. Details of Our Graph Query Translation Prompt
As described in Section 3.2.2, we used large language models to translate the candidate graph queries to natural language. Specifically, our prompt uses the following structure:
We defined a Graph Query to locate target data on mobile UI structure. It describes the unique attributes of the data we are targeting. For example, the query
(conj (HAS\_CLASS\_NAME android.widget.FrameLayout) (RIGHT (conj (hasText 6) (HAS\_CLASS\_NAME android.widget.TextView) (HAS\_PACKAGE\_NAME com.ubercab)) ) (HAS\_PACKAGE\_NAME com.ubercab))
stands for: the information that is located to the right of a text “6”
Below I have a few queries, can you help me translate them to human-readable format like above?
  1. (1)
    …(first query)
  2. (2)
    …(second query)
  3. (3)
    …(etc.)
Leave out UI element names (TextView, FrameLayout) and do not make any reference to “view”. Users only care about the data and information contained in the view instead of the UI elements themselves. “With numeric index xx” should be translated into “the xx in the list” (first, second, etc.). Be as concise as possible. Make sure you return the translation in the order I presented the queries above, separated by new lines. Return nothing else.
A.2. Evaluation Study 1 Participant Demographics
Here we present the demographics of our evaluation Study 1 participants, as discussed in Section 4.2.
ID Age Gender Research Experience Collected Data Types Data Collection Tools Data Collection Projects
P1 23 Male 3 years IDE logs, eye tracking, semi-structured interview, survey IDE plugin, screen recorder, audio recorder, eye tracker 2
P2 30 Female 5 years app usage, surveys, interviews Google Analytics, Amazon Web Services, think-aloud, interviews, surveys, Qualtrics, MAXQDA About 6 (5–10)
P3 26 Male 4 UI interaction, think-aloud, interview Computer program, recorder 4
P4 32 Female 2 years Interviews Audio recording 2
P5 26 Female 3.5 years Gaze, image description Eye tracker 2
Table 4. The demographics of our evaluation Study 1 participants, as discussed in Section 4.2.
[…]
This table summarizes demographic and research background information for five participants in Evaluation Study 1. Each row represents a participant (P1 to P5), and columns include age (ranging from 23 to 32), gender (three female, two male), and years of research experience (from 2 to 5 years). The table also lists the types of data each participant has collected in past projects, such as IDE logs, eye tracking, interviews, surveys, and gaze data. Tools used for data collection include screen recorders, audio recorders, IDE plugins, eye trackers, and online survey platforms. The final column reports the number of data collection projects each participant has worked on, ranging from 2 to about 6.
A.3. Evaluation Study 2 Participant Demographics
Here we present the demographics of our evaluation Study 1 participants, as discussed in Section 4.3.
ID Age Gender Instagram Frequency Uber Frequency Chrome Mobile Frequency
PB1 21 Male Every week Once in a few weeks Everyday
PB2 25 Male Everyday Once in a few months Everyday
PB3 23 Male Every week Every week Everyday
PB4 30 Male Everyday Once in a few weeks Everyday
PB5 26 Female Everyday Once in a few weeks Everyday
Table 5. Demographics information and app usage frequency for our evaluation Study 2 participants in Section 4.3.
[…]
This table presents demographic and app usage frequency data for five participants in Evaluation Study 2. Each row corresponds to a participant (PB1 to PB5), with columns for age (from 21 to 30), gender (three male, two female), and self-reported frequency of using three mobile apps: Instagram, Uber, and Chrome. Instagram usage ranges from “Every week” to “Everyday,” Uber usage varies from “Every week” to “Once in a few months,” and all participants reported using Chrome on mobile “Everyday.” The table provides context on participants’ familiarity with the apps used during the study.